"A water sample contains following impurities: Mg(HCO3)2 = 292 mg/L, CaCl2 = 166.5 ppm, SiO2 = 5 ppm, Calculate the volume of EDTA solution (0.1 N) required to titrate 50 mL of the water sample before boiling and after boiling-filtering. (Ans: 1.5 mL)" 


1
Expert's answer
2020-07-14T15:18:00-0400

Solution.

"m(Mg(HCO3)2) = 292 \times 0.05 = 14.6 \ mg"

"m(CaCl2) = 166.5 \times 0.05 = 8.325 \ mg"

"\nu = \nu(Mg(HCO3)2) + \nu(CaCl2) = \frac{0.0146}{146} + \frac{0.008325}{111} = 0.000175 \ mol"

"V(EDTA) = \frac{\nu \times 1000}{H} = 1.75 \ ml"

Answer:

V(EDTA) = 1.75 ml